# Bohr-Sommerfeld type surgeries and Disk Potential

### v1

@@ 3.2. Convention for first homology
We now select a splitting of $H_{1}(BSP_{0}(L))$ corresponding to the splitting in A.2 which we use to state the change in disk-potential formula. The $\zeta$ component in the splitting A.2 is of main importance. By a slight abuse of notation, let $\zeta:S^{1}\rightarrow L$ be the cycle representing $\zeta$ such that it intersects $i(\Lambda\times\{1\})$ at $p$, say $\zeta(\theta_{1})=p.$ Note
(3.2.1) $H_{1}(BSP_{0}(L))=i_{*}(\Lambda\times[0,1])\oplus\mathbb{Z}\langle\widetilde{\zeta}\rangle\oplus R,$
@@
We will choose the path $\psi(t)$ to be equal to $l(1-t)$ where $l$ is the lift of the boundary $\partial u,$ of a Maslov 2 disk $u$ which has non-zero contribution in the disk potential $W_{L_{Y}}$.
@@ 3.3. Main theorem
We can now state the main theorem of this section,
@@
(3.3.1) $W_{BSP_{0}(L)}(x_{1},\dots,x_{k},z,w_{1},\dots,w_{l})=W_{L}(x_{1},\dots,x_{k},zW_{\Lambda}(x_{1},\dots,x_{k}),w_{1},\dots,w_{k}),$
@@
When we perform a zero-BSP surgery with profile $\Lambda=S^{1}=S^{3}\cap\mathbb{R}^{2}\hookrightarrow S^{3}$ where we view $S^{3}$ as a unit sphere in $\mathbb{C}^{2}$, $Y=\mathbb{C}P^{1}$, $L_{Y}=\mathbb{R}P^{1}$, our wall-crossing formula matches those which are already in the literature, see [Aur07, PW19, ch10]. Indeed, notice that performing such a zero area BSP is the same as performing a Lagrangian disk-surgery and the change of potential is given by
$W_{BSP_{0}(L)}(x,y,w_{1},...)=W_{L}(x,y(1+x),w_{1},...),$

### v2

@@ 3.2. Convention for first homology
We select a splitting of $H_{1}(BSP_{0}(L))$ corresponding to the splitting in A.2 which will be used to state the change in disk-potential formula. The $\zeta$ component in the splitting A.2 is of main importance. By a slight abuse of notation, let $\zeta:S^{1}\rightarrow L$ be the cycle representing $\zeta$ such that it intersects $i(\Lambda\times\{1\})$ at $p$, say $\zeta(\theta_{p})=p.$ Note
(3.2.1) $H_{1}(BSP_{0}(L))=i_{*}(H_{1}(\Lambda\times\{0\})\oplus\mathbb{Z}\langle\widetilde{\zeta}\rangle\oplus R,$
@@
We will choose the path $\psi(t)$ to be equal to l(1t)l(1-t) where $l$ is the lift of the boundary $\partial u,$ of a Maslov 2 disk $u$ which has non-zero contribution in the disk potential $W_{L_{Y}}$.
@@ 3.3. Main theorem
We can now state the main theorem of the article, which is a precise version of Theorem 1.
@@
(3.3.1) $W_{BSP_{0}(L)}(x_{1},\dots,x_{k},z,w_{1},\dots,w_{l})=W_{L}(x_{1},\dots,x_{k},zW_{\Lambda}(x_{1},\dots,x_{k}),w_{1},\dots,w_{l}),$
@@
When we perform a zero-BSP surgery with profile $\Lambda=S^{1}=S^{3}\cap\mathbb{R}^{2}\hookrightarrow S^{3}$ where we view $S^{3}$ as a unit sphere in $\mathbb{C}^{2}$, $Y=\mathbb{C}P^{1}$, $L_{Y}=\mathbb{R}P^{1}$, our wall-crossing formula matches those which are already in the literature, see [Aur07, PW19, ch10]. Indeed, notice that performing such a zero area BSP is equivalent to performing a Lagrangian disk-surgery and the change of potential is given by
[…]